A regulation requires that in order for a soda to be termed as a "naturally flavored," the drink must contain at least 10% fruit juice. How much pure juice must a manufacturer add to 900 gallons of orange soda to conform to the regulation?
Answer by richwmiller(17219)

x/(900+x)=1/10
10x=900+x
9x=900
x=100